find the width of a rectangle if area is 8 3/4 and length is 3 1/2 show your work
Answer by Cromlix(4381)

Area of rectangle = Length * Width
8 3/4 = 3 1/2 * Width
Width = 8 3/4/ 3 1/2
8 3/4 = 35/4
3 1/2 = 7/2
35/4 divided by 7/2
(35 * 2)/(4 * 7)
= 5/2 = 2 1/2
Width = 2 1/2 units.
Hope this helps.
:-)